OpenStudy (anonymous):

If h(x) = −5x − 10, find h(−2).

OpenStudy (hoblos):

put (-2) in the place of x h(−2)= -5(-2) -10 = 10-10 =0

OpenStudy (anonymous):

Here , h is a function of x , so putting x=-2 , will make h a function of -2 and will give its value: h(-2)=-5(-2)-10=10-10=0

OpenStudy (muhammad_nauman_umair):

h(-2)=0